Let f(x,y)={y^3/x^2+y^2 , 0 otherwise if (x,y) is not =(0,0) show that f is continuous but not differentiable at (0,0)

1
Expert's answer
2022-05-05T07:27:10-0400

Since it is not clear how the function looks like, consider two cases:

  1. The function "f(x,y)" has the form: "f(x,y)=\frac{y^3}{x^2}+y^2". Set "y=t", "x=t^2". We receive "f(t^2,t)=\frac{t^3}{t^4}+t^2" . This function is not continuous at "(0,0)", since it approaches "\infty" as "(x,y)\rightarrow(0,0)" along the curve "x=t^2,y=t". Thus, it is not differentiable at "(0,0)."
  2. The function "f(x,y)" is: "f(x,y)=\frac{y^3}{x^2+y^2}". Consider polar coordinates: "x=r\cos\phi", "y=r\sin\phi". We get: "f(r\cos\phi,r\sin\phi)=\frac{r^3\sin^3\phi}{r^2}=r\sin^3\phi". The latter approaches as "(r,\phi)\rightarrow(0,0)". It corresponds to "(x,y)\rightarrow(0,0)". Thus, we can put "f(0,0)=0" and the function will be continuous at "(0,0)". By definition, the function "f" is differentiable at "(0,0)", in case there are numbers "a" and "b" such that "\lim_{(x,y)\rightarrow(0,0)}\frac{f(x,y)-ax-by}{\sqrt{x^2+y^2}}=0".

Consider polar coordinates: "x=r\cos\phi,y=r\sin\phi". Then, "\frac{f(x,y)-ax-by}{\sqrt{x^2+y^2}}=\sin^3\phi-a\cos\phi-b\sin\phi". Thus, for "a=0" we receive: "\lim_{(x,y)\rightarrow(0,0)}\frac{f(x,y)-ax-by}{\sqrt{x^2+y^2}}=0". Therefore, the function is differentiable at "(0,0)."